OpenStudy (anonymous):

Ignore air resistance. A diver starts 30 feet above the water (about the height of an Olympic platform diver). What is the diver’s velocity at impact?

OpenStudy (anonymous):

@mukushla can u help?

OpenStudy (anonymous):

no its calculus

OpenStudy (anonymous):

soo is it correct?

OpenStudy (anonymous):

okay :/

OpenStudy (anonymous):

@rulnick @rizwan_uet @RadEn can u help?

OpenStudy (anonymous):

use second equation of motion 2as = vf^2 -vi^2

OpenStudy (anonymous):

put a = g =9.8 s = 30 feet = ........meters do conversion vi =0 now you can find vf

OpenStudy (anonymous):

okay thank YOU! :)

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

ok so its like this? 2(9.8) (9.14) = vf^2 - 0^2

OpenStudy (anonymous):

@dumbcow can u check this?

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

hi

OpenStudy (dan815):

how long does it take him to fall down 30 feet?

OpenStudy (dan815):

when you get that time multiply it by acceleration as accel=velocity/time so velocity=accel*time time you can get by solving this equation for t h=1/2at^2 where h = displacement